# Curriculum learning based pre-training using Multi-Modal Contrastive Masked Autoencoders

### v1

@@ 3.4 Denoising
Denoising would encourage the encoder to extract high frequency components which can be complementary to the MAE pre-training in the second stage of our curriculum. The computation of the denoising doesn’t bring any overhead to the pre-training compared to the vanilla MAE, since we use a lightweight MLP and compute the loss on the reconstructed unmasked patches. The decoder also reconstructs the unmasked patches which may go wasted in the vanilla MAE pre-training. We also provide an ablation study (c.f. Table 6) on the effect of each component in the denoising stage. Lastly, the figure 3 shows the overall setup of the denoising.
3.5 Pre-training Curriculum
We are now ready to present our full curriculum for pre-training ViT using RGB-D datasets. In the first stage, we pre-train the model using contrastive loss as mentioned in section 3.2.
@@ 3.5 Pre-training Curriculum
In the second stage, we pre-train the model using Masked autoencoding and denoising. Therefore, the overall stage-2 objective can be defined as:
$\mathcal{L}_{\mathrm{stage2}}=\alpha\mathcal{L}_{\mathrm{depth}}+\beta\mathcal{L}_{\mathrm{denoise}}$ (5)
Where $\alpha$, $\beta$ are the weights for each loss function during pre-training. Please see appendix on the details of these hyper-parameters.
@@ 4 Results
In this section, we show the efficacy of our approach for ViT-B backbone on semantic segmentation, instance segmentation and depth estimation. We pre-train ViT-B on the ScanNet [22] dataset which contains 2.5M RGB-D frames from 1513 training video sequences. We regularly sample every 25th frame without any filtering during pre-training following Mask3D [43]. To study the effectiveness of our pre-training approach for small scale datasets, we pre-train the model using SUN RGB-D dataset [77]. It contains 10,335 RGB-D images captured from 4 different sensors. We use the official split of train/test for pre-training and evaluation. Furthermore, we initialize our encoders using ImageNet self-supervised weights [38]. Please refer to the appendix for more details on implementations and hyper-parameters.
@@
Mask3D [43] Depth ViT-B ImageNet+ScanNet RGB 66.0
@@
Ours Depth ViT-B ImageNet+ScanNet RGB 67.0
@@
Table 1: ScanNet Semantic Segmentation. Our approach outperforms Mask3D and other SOTA approaches that leverage RGB-D data during pre-training. * means MultiMAE was pre-trained using ScanNet by initializing the ViT-B encoder with ImageNet SSL weights.
@@ 4.1 Semantic Segmentation
We first present the results for semantic segmentation. Following MultiMAE [6], we use ConvNext architecture [59] as a segmentation decoder for our approach and the baselines. For evaluation, we use mean Intersection over Union (mIoU). Table 1 shows the performance of our approach and other baselines on ScanNet 2D semantic segmentation task. For fine-tuning, we follow Mask3D [43] and sample every 100th frame, resulting in 20,000 training images and 5000 validation images. Our approach achieves 67.0 mIoU compared to 66.0 mIoU for Mask3D and 65.1 mIoU for MultiMAE. For fair comparison, we pre-train MultiMAE with RGB and depth only. Moreover, compared to MAE pre-trained on ScanNet RGB modality, our approach improves the performance by +2.5 mIoU. Furthermore, compared to Pri3D, a multi-view contrastive learning approach, we outperforms it by +6.8 mIoU. Table 2 shows the results on SUN RGB-D 2D semantic segmentation. We observe that our method outperforms other state-of-the-art approaches. More notably, we achieve 48.1 mIoU compared to 47.4 mIoU for Mask3D and 47.1 mIoU for MultiMAE. These results demonstrate that our pre-training approach can be used for relatively small scale RGB-D datasets. Please follow supplementary materials for more results on NYUv2 semantic segmentation.
@@
Ours Depth ViT-B ImageNet+SUN RGB-D RGB 48.1
@@
4.2 Depth Estimation
In this section, we study the efficacy of our pre-trained model for depth estimation task. We use NYUv2 [63] for this downstream task and report $\delta_{1}$ on the test set which shows the percentage of pixels that have an error ratio ($\max\{\frac{\hat{y}_{p}}{y_{p}},\frac{y_{p}}{\hat{y}_{p}}\}$) below 1.25 [25]. We use DPT [71] as dense prediction head following MultiMAE for all the approaches and use the model pre-trained on ScanNet dataset for fine-tuning. The results on the depth prediction task are reported in Table 3. We can draw the same observation from the table that our approach outperforms state-of-the-art approaches. Specifically, it achieves 86.8 $\delta_{1}$ compared to 85.4 $\delta_{1}$ for Mask3D and 85.3 $\delta_{1}$ for MultiMAE. We also compare our approach with CroCo [85], a cross-view completion pre-training approach specifically designed for 3D vision tasks. In particular, we achieve an improvement of +1.2 $\delta_{1}$ over CroCo. These results again verify the generalizibility of our curriculum based pre-training strategy by transferring the representations to various downstream tasks and datasets.
@@ 4.2 Depth Estimation
Ours Depth ViT-B ImageNet+ScanNet RGB 86.8
@@
4.3 Instance Segmentation
@@
Ours ImageNet + ScanNet 23.3
@@
Curriculum order mIoU
Contrastive $\rightarrow$ MIM + denoising 67.0
MIM + denoising $\rightarrow$ Contrastive 64.9
Contrastive + MIM + denoising 64.2
Table 5: Effect of ordering of pre-training strategies on ScanNet 2D semantic segmentation.
@@ 4.4 Ablation Studies
Effect of curriculum order.
We study the impact of changing the order of learning in our curriculum. Table 5 shows the performance on ScanNet 2D semantic segmentation. It can be seen that our proposed curriculum strategy outperforms the other ones. More specifically, when we pre-train the model in a single framework, it performs the worst overall. We hypothesize that masking the inputs degrades the optimization of contrastive learning as it is important to learn pixel-to-pixel level or patch-level cross-modal relationships. Moreover, by pre-training using the masked image modeling first and then contrastive learning, we achieve 64.9 mIoU compared to 67.0 mIoU. This shows the importance of learning the easier task first.
Ablation study: Denoising
We consider three settings in this ablation study: 1) Without any noise 2) Input with noise only 3) full method with all the denoising components. Table 6 shows the performance of the different components of denoising loss on ScanNet 2D segmentation. It can be seen from the table that by simply adding noise in the input, the performance improves 0.5 mIoU. It is further improved by 0.8 mIoU with full denoising method that consists of noise prediction and positional encoding in the decoder.
Ablation study: Loss functions.
In Table 8, we highlight the importance of using different self-supervised frameworks in a curriculum manner. If we drop contrastive learning, our approach would only consists of Masked Image modeling and denoising, leading to a performance drop to 66.3 mIoU highlighting the importance of having constrastive learning in the first stage. Nonetheless, it is still marginally better than Mask3D (66.0 mIoU) implying that the learned representations using MIM and noise prediction are more discriminative and better transferable to the downstream task.
@@
w/out noise ViT-B 66.2
noise only ViT-B 66.7
Full ViT-B 67.0
@@
MIM + Denoising ViT-B 66.3
Depth + RGB Reconstruction ViT-B 66.4
Full ViT-B 67.0
Table 8: Ablation study of different loss functions on ScanNet 2D semantic segmentation.
@@
We also study the effect of reconstructing the RGB input besides the depth modality. It can be seen from the table 8 that it doesn’t improve the performance. One plausible reason for this is that reconstructing the RGB modality is a simpler task during pre-training, given the MAE-based initialization and 3D priors are also encoded during the depth prediction.
Results on different ViT variants.
@@
Performance without ImageNet pre-training.
In Table 10, we show the performance on ScanNet 2D segmentation without initializing the models with ImageNet weights before pre-training. For this setting, we use ConvNext based decoder across all the approaches. We see a significant drop in the performance without ImageNet based initialization. Nevertheless, our approach still outperform Mask3D by +0.8 mIoU.
Method Backbone mIoU
Train from Scratch ViT-B 32.6
Mask3D [43] ViT-B 41.3
Ours ViT-B 42.1
Table 10: Without ImageNet initialization, our approach still outperforms Mask3D.
@@
4.5 Low-data Regime Scenarios
We show the data-efficient nature of our pre-training strategy by fine-tuning it in limited labeled data scenarios. Figure 4 shows the performance on ScanNet 2D semantic segmentation. Our approach consistently outperform Mask3D, MAE and Pri3D across different ranges of limited labeled data. More notably, it almost achieves the same performance as MAE with only 60% of the original training set.
@@ 5 Conclusion
We propose a new curriculum learning based pre-training strategy that effectively combines powerful self-supervised learning methods in a single framework to pre-train ViTs using RGB-D datasets. Our curriculum consists of two phases of learning: (1) contrastive learning which aligns the RGB and depth patches to extract discriminative local context features, (2) Masked Image Modeling which reconstructs the masked patches in the depth modality using RGB-D input to learn spatial statistical correlations. To learn high frequency features, we introduce a denoising loss inspired by the recent success of diffusion models to predict the noise added in the input. We demonstrate the effectiveness of our approach on various downstream tasks such as semantic segmentation, depth prediction and instance segmentation. Finally, we show scalability and data-efficiency by pre-training the models using small scale RGB-D datasets.

### v2

@@ 3.4 Denoising
Denoising would encourage the encoder to extract high frequency components which can be complementary to the MAE pre-training in the second stage. The computation of the denoising doesn’t bring any overhead to the pre-training compared to the vanilla MAE, since we use a lightweight MLP and compute the loss on the reconstructed unmasked patches. The decoder also reconstructs the unmasked patches which may go wasted in the vanilla MAE pre-training. We also provide an ablation study (c.f. Table 6) on the effect of each component in the denoising stage. Lastly, the figure 2 shows the overall setup of the denoising.
3.5 Feature Distillation
To distill the global modality-specific embeddings in the second stage, we use the output of the encoders as the visual embeddings for each modality. We then apply max pooling to output a global embeddings. Afterwards, we align this second-stage global embedding to the global visual embedding from the first-stage model using smooth $\ell_{1}$ loss which is given as:
$\mathcal{L}_{\mathrm{distill}}(\mathbf{f_{2}},\mathbf{f_{1}})=\begin{cases}\frac{1}{2}(\mathbf{f_{2}}-\mathbf{f_{1}})^{2}/\beta,&|\mathbf{f_{2}}-\mathbf{f_{1}}|\leq\beta\\ (|\mathbf{f_{2}}-\mathbf{f_{1}}|-\frac{1}{2}\beta),&\text{otherwise}\end{cases},$ (4)
where $\beta$ is set to 2.0; $\mathbf{f_{2}}$ and $\mathbf{f_{1}}$ are the max-pooled features from the stage-2 and stage-1 encoders respectively. In practice, we employ the feature distillation for both RGB and depth. So, the final distillation loss is the sum of modality-specific distillation losses.
3.6 Progressive Pre-training
We are now ready to present our full algorithm for pre-training ViT using RGB-D datasets. In the first stage, we pre-train the model using contrastive loss as mentioned in section 3.2.
@@ 3.6 Progressive Pre-training
In the second stage, we pre-train the model using Masked autoencoding, denoising and feature distillation. Therefore, the overall stage-2 objective can be defined as:
$\mathcal{L}_{\mathrm{stage2}}=\alpha\mathcal{L}_{\mathrm{depth}}+\beta\mathcal{L}_{\mathrm{denoise}}+\gamma\mathcal{L}_{\mathrm{distill}}$ (6)
Where $\alpha$, $\beta$ and $\gamma$ are the weights for each loss function during pre-training. Please see appendix on the details of these hyper-parameters.
@@ 4 Results
In this section, we show the efficacy of our approach for ViT-B backbone on semantic segmentation, instance segmentation and depth estimation. In the first stage, we pre-train the ViT-B encoders using ImageNet [22] dataset. We use off the shelf depth estimation model [29] to extract the depths. In the second stage, we continue pre-training the encoders on the ScanNet [21] dataset which contains 2.5M RGB-D frames from 1513 training video sequences. We regularly sample every 25th frame without any filtering during pre-training following Mask3D [43]. To study the effectiveness of our pre-training approach for small scale datasets, we also pre-train the model using SUN RGB-D dataset [73] in the second stage. It contains 10,335 RGB-D images captured from 4 different sensors. We use the official split of train/test for pre-training and evaluation. We also initialize the second stage encoders with the weights learnt in the first stage. Please refer to the appendix for more details on implementations and hyper-parameters.
@@
Mask3D** [43] Depth ViT-B ImageNet+ScanNet RGB 66.2
@@
Ours Depth ViT-B ImageNet+ScanNet RGB 67.5
@@
Table 1: ScanNet Semantic Segmentation. Our approach outperforms Mask3D and other SOTA approaches that leverage RGB-D data during pre-training. * means MultiMAE was pre-trained using ScanNet by initializing the ViT-B encoder with ImageNet SSL weights. ** means ConvNext decoder is used instead of transformer decoder.
@@ 4.1 Semantic Segmentation
We first present the results for semantic segmentation. Following MultiMAE [6], we use ConvNext architecture [57] as a segmentation decoder for our approach and the baselines. All the baselines are initialized with ImageNet self-supervised weights [37]. For evaluation, we use mean Intersection over Union (mIoU). Table 1 shows the performance of our approach and other baselines on ScanNet 2D semantic segmentation task. For fine-tuning, we follow Mask3D [43] and sample every 100th frame, resulting in 20,000 training images and 5000 validation images. Our approach achieves 67.5 mIoU compared to 66.2 mIoU for Mask3D and 65.1 mIoU for MultiMAE. For fair comparison, we pre-train MultiMAE with RGB and depth only and use ConvNext decoder for Mask3D. Moreover, compared to MAE pre-trained on ScanNet RGB modality, our approach improves the performance by +3.0 mIoU. Furthermore, compared to Pri3D, a multi-view contrastive learning approach, we outperforms it by +7.3 mIoU. Table 2 shows the results on SUN RGB-D 2D semantic segmentation. We observe that our method outperforms other state-of-the-art approaches. More notably, we achieve 48.7 mIoU compared to 47.4 mIoU for Mask3D and 47.1 mIoU for MultiMAE. Table 4 shows the comparison of our approach and CoMAE [83] on the SUN RGB-D dataset in terms of mean accuracy, as CoMAE did not report mIoU results for this dataset. We observe that our approach outperforms the CoMAE by 1.1%. These results demonstrate that our second-stage pre-training is effective even on relatively small-scale RGB-D datasets. We also present results on NYUv2 semantic segmentation, where similar observations can be made when comparing our method with the competing approaches. For further details, please refer to the supplementary materials.
@@
Ours Depth ViT-B ImageNet+SUN RGB-D RGB 48.7
@@
Ours Depth ViT-B ImageNet+ScanNet RGB 87.1
@@
Methods Pre-train accs
CoMAE [83] SUN RGB-D 64.3
Ours ImageNet + SUN RGB-D 65.4
Table 4: Comparison with CoMAE [83] using average accuracy (accs) on SUN RGB-D Semantic Segmentation.
4.2 Depth Estimation
In this section, we study the efficacy of our pre-trained model for depth estimation task. We use NYUv2 [61] for this downstream task and report $\delta_{1}$ on the test set which shows the percentage of pixels that have an error ratio ($\max\{\frac{\hat{y}_{p}}{y_{p}},\frac{y_{p}}{\hat{y}_{p}}\}$) below 1.25 [25]. We use DPT [69] as dense prediction head following MultiMAE for all the approaches and use the model pre-trained on ScanNet dataset for fine-tuning. The results on the depth prediction task are reported in Table 3. We can draw the same observation from the table that our approach outperforms state-of-the-art approaches. Specifically, it achieves 87.1 $\delta_{1}$ compared to 85.4 $\delta_{1}$ for Mask3D and 85.3 $\delta_{1}$ for MultiMAE. We also compare our approach with CroCo [79], a cross-view completion pre-training approach specifically designed for 3D vision tasks. In particular, we achieve an improvement of +1.5 $\delta_{1}$ over CroCo. These results again verify the generalizibility of our two-stage pre-training strategy by transferring the representations to various downstream tasks and datasets.
@@ 4.2 Depth Estimation
Ours ImageNet + ScanNet 23.7
@@
4.3 Instance Segmentation
@@ 4.4 Ablation Studies
w/out noise ViT-B 66.5
noise only ViT-B 66.9
Full ViT-B 67.5
@@
Denoising.
We consider three settings in this ablation study: 1) Without any noise 2) Input with noise only 3) full method with all the denoising components. Table 6 shows the performance of the different components of denoising loss on ScanNet 2D segmentation. It can be seen from the table that by simply adding noise in the input, the performance improves 0.4 mIoU. It is further improved by 0.6 mIoU with full denoising method that consists of noise prediction and positional encoding in the decoder.
Contrastive Reconstruction Denoising Distill mIoU
63.4
66.3
66.5
67.0
67.5
Table 7: Effect of different pre-training objectives on ScanNet 2D semantic segmentation.
Effect of pre-training objectives.
In Table 7, we highlight the importance of using different self-supervised frameworks. It shows that contrastive learning alone can perform reasonably well. By pre-training using only depth reconstruction loss in the second stage, the performance jumps from 63.4 to 66.3 which is on-par with Mask3D while integrating all three objectives in the stage-2 significantly enhances the final performance. Lastly, without our feature distillation, the performance drops from 67.5 to 67.0 mIoU.
@@
Depth + RGB Reconstruction ViT-B 66.9
Full ViT-B 67.5
Table 9: Ablation study of RGB and Depth reconstruction in the second stage on ScanNet 2D semantic segmentation.
@@
We also study the effect of reconstructing the RGB input besides the depth modality. It can be seen from the table 9 that it doesn’t improve the performance. A plausible explanation for this is that reconstructing the RGB modality during pre-training is relatively simpler, as the 3D priors are encoded through the depth prediction process, providing additional reinforcement.
@@
Results on different ViT variants.
4.5 Low-data Regime Scenarios
We show the data-efficient nature of our pre-training strategy by fine-tuning it in limited labeled data scenarios. Figure 3 shows the performance on ScanNet 2D semantic segmentation. Our approach consistently outperform Mask3D, MAE and Pri3D across different ranges of limited labeled data. More notably, it marginally outperforms MAE with only 60% of the original training set.
@@ 5 Conclusion
We propose a new progressive pre-training strategy that effectively combines powerful self-supervised learning methods in a single framework to pre-train ViTs using RGB-D datasets. Our pre-training consists of two phases of learning: (1) contrastive learning which aligns the RGB and depth patches to extract discriminative local context features, (2) Masked Image Modeling which reconstructs the masked patches in the depth modality using RGB-D input to learn spatial statistical correlations and embed 3D priors. To learn high frequency features, we introduce a denoising loss inspired by the recent success of diffusion models to predict the noise added in the input. To further leverage the knowledge learnt in the stage-1, we employ feature distillation based on the global embeddings. We demonstrate the effectiveness of our approach on various downstream tasks such as semantic segmentation, depth prediction and instance segmentation. Finally, we show scalability and data-efficiency by pre-training the second stage using small scale RGB-D datasets.